# Double Robustness of Local Projections

### v2

@@ 3.1 Large-sample distributions and confidence interval coverage
See Section B.2. ∎
The convergence rate $T^{-\min\{1/2,\zeta\}}$ of the VAR estimator is potentially slower than the $T^{-1/2}$ rate achieved by LP. This is because the VAR estimator suffers from bias of order $T^{-\zeta}$, while the stochastic terms of order $T^{-1/2}$ are the same as they would be in a correctly specified SVAR($p$) model. [^3] The first stochastic term captures sampling uncertainty in the reduced-form impulse responses $\hat{A}^{h}$, while the second term captures uncertainty in the structural impact response vector $\hat{\nu}$. The VAR bias is only asymptotically negligible if $\zeta>1/2$, a much smaller degree of robustness than shown above for LP. The case $\zeta=1/2$ is of particular interest, as then the bias and standard deviation are of the same asymptotic order (see also Schorfheide, 2005). MA terms of order $T^{-1/2}$ can be detected with asymptotic probability strictly between 0 and 1 by specification tests, as will be shown in Section 3.2.
The asymptotic bias is due to two forces: first, the coefficient matrix $\hat{A}$ is biased due to the endogeneity caused by the MA terms, and second, the VAR estimator extrapolates the $h$ impulse response based on a parametric formula $\hat{A}^{h}$ that does not hold exactly in the true VARMA model (2.1). This is more easily seen in the special case of a univariate model $y_{t}=\rho y_{t-1}+[1+T^{-\zeta}\alpha(L)]\varepsilon_{t}$ with $n=m=1$, in which case
@@
where $\hat{\rho}=\hat{A}$ is the AR(1) coefficient from an OLS regression of $y_{t}$ on $y_{t-1}$. [^4] Lag augmentation of the VAR impulse response estimator as in Inoue and Kilian (2020) may reduce the first term in the bias formula, but it does not affect the second term.
@@
The preceding results imply that the conventional LP confidence interval is robust to misspecification while the conventional VAR interval is not. We define the level-($1-a$) LP and VAR confidence intervals using the standard formulae:
@@
Here $z_{1-a/2}$ is the $21-a/$ 2 quantile of the standard normal distribution, and $\operatorname{aVar}(\hat{\beta}_{h})$ and $\operatorname{aVar}(\hat{\delta}_{h})$ are the asymptotic variances of the leading (order-$T^{-1/2}$) stochastic terms in the representations of the LP and VAR estimators in Propositions 3.1 and 3.2; explicit formulae for the asymptotic variances are given in Corollary A.2 in Section A.3, which also implies that $\operatorname{aVar}(\hat{\beta}_{h})\geq\operatorname{aVar}(\hat{\delta}_{h})$. None of the results below would change if we replaced the asymptotic variances with the conventional consistent estimates of these (that assume correct specification, as implemented in standard econometric software packages). [^5] Under Assumption 2.1, homoskedastic OLS standard errors suffice. In practice we recommend heteroskedasticity-robust standard errors or the bootstrap, as in Montiel Olea and Plagborg-Møller (2021).
@@
Under Assumption 2.1, $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=1-a$. If moreover $\operatorname{aVar}(\hat{\delta}_{h})>0$ and $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$, then $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\delta}_{h}))=\lim_{T\to\infty}\{1-r\left(T^{1/2-\zeta}b_{h};z_{1-a/2}\right)\}$, where $b_{h}\equiv\operatorname{aBias}(\hat{\delta}_{h})/\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}$, $r(b;c)\equiv P_{Z\sim N(0,1)}(|Z+b|>c)=\Phi(-c-b)+\Phi(-c+b)$, and $\Phi(\cdot)$ is the standard normal distribution function.
@@
Considering separately the three cases $\zeta\in(1/4,1/2)$, $\zeta=1/2$, and $\zeta>1/2$, the result is an immediate consequence of Propositions 3.1 and 3.2. ∎
LP robustly controls coverage, while the VAR confidence interval generically has coverage converging to zero for $\zeta\in(1/4,1/2)$, and strictly below the nominal level $1-a$ for $\zeta=1/2$. Intuitively, the VAR confidence interval has the right width (the same as in the correctly specified case) but the wrong location due to the bias, thus causing coverage distortions.
@@ 3.2 Hausman misspecification test
Impose Assumption 2.1 and assume $\operatorname{aVar}(\hat{\beta}_{h})>\operatorname{aVar}(\hat{\delta}_{h})>0$. Then the asymptotic rejection probability of the Hausman test equals
@@
where $b_{h}$ and $r(\cdot,\cdot)$ were defined in Corollary 3.1.
@@
Considering separately the three cases $\zeta\in(1/4,1/2)$, $\zeta=1/2$, and $\zeta>1/2$, the result is an immediate consequence of Propositions 3.1 and 3.2 as well as Corollary A.2 in Section A.3. ∎
As claimed previously, the Hausman test is consistent against MA misspecification of order $T^{-\zeta}$ with $\zeta\in(1/4,1/2)$, except in the knife-edge case where $\operatorname{aBias}(\hat{\delta}_{h})=0$. When $\zeta=1/2$ and $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$, the asymptotic rejection probability is strictly between the significance level $a$ and 1. In Section 4 we will use the Hausman test to quantify the difficulty of detecting especially pernicious types of model misspecification.
3.3 The role of lag length
One simple way to remove the asymptotic bias of the VAR estimator is to control for sufficiently many lags, since in this case the estimator is asymptotically equivalent with the LP estimator. The larger the impulse horizon of interest, the more lags are required for bias reduction. See Plagborg-Møller and Wolf (2021) and Xu (2023) for related results in models without explicit MA misspecification.
@@ 3.3 The role of lag length
Suppose the model (2.2) written in companion form (2.1) satisfies Assumption 2.1. Let $\tilde{\check{y}}_{t}$ denote the stationary solution to equation (2.2) when $\alpha(L)=0$. If $\varepsilon_{j^{*},t-\ell}\in\operatorname{span}(\tilde{\check{y}}_{t-1},\dots,\tilde{\check{y}}_{t-p})$ for all $\ell=1,\dots,h$, then $\operatorname{aBias}(\hat{\delta}_{h})=0$ and $\operatorname{aVar}(\hat{\delta}_{h})=\operatorname{aVar}(\hat{\beta}_{h})$. In particular, these results obtain if either of the following two sufficient conditions hold:
@@
See Section B.4. ∎
We show in Section 4 that controlling for so many lags that LP and VAR are asymptotically equivalent is in fact the only way to guarantee that the asymptotic bias of the VAR estimator is zero.
4 Some unpleasant VARithmetic
To show that the fragility of VARs is likely to matter in practice, we now investigate the worst-case properties of VAR procedures under a tight constraint on the amount of misspecification. We prove that there is no free lunch: the conventional VAR confidence interval is robust to misspecification if, and only if, the LP and VAR intervals coincide asymptotically. VARs with short-to-moderate lag lengths instead suffer from severe coverage distortions for small amounts of misspecification that is hard to rule out economically or statistically. Beyond increasing the lag length, an alternative strategy to fix VAR undercoverage is to use a larger bias-aware critical value; however, we show that the resulting confidence interval is usually wider than the LP interval. Finally, we generalize all our results to the case of joint inference on multiple impulse responses.
@@ 4.1 Worst-case bias and mean-squared error
where we define the norm
$\|\alpha(L)\|\equiv\sqrt{\sum_{\ell=1}^{\infty}\operatorname*{trace}\{D\alpha_{\ell}^{\prime}D^{-1}\alpha_{\ell}\}}.$
Suppose we are willing to impose a priori that the noise-to-signal ratio is at most $M^{2}/T$ for some constant $M\in(0,\infty)$. For small $M^{2}/T$, this roughly means that a fraction $M^{2}/T$ of the variance of the model’s error term is due to the misspecification. This corresponds to restricting the parameter space for $\alpha(L)$ to all absolutely summable lag polynomials that satisfy $\|\alpha(L)\|\leq M$. In the following we will consider the worst-case properties of the VAR estimator over this parameter space, treating the other (consistently estimable) parameters $(A,H,D)$ as fixed.
@@
Impose Assumption 2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\delta}_{h})>0$. Then
@@
where we recall the definition $b_{h}=\operatorname{aBias}(\hat{\delta}_{h})/\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}$.
@@
The claim is a special case of Proposition 4.2 below. ∎
@@
Proposition 4.1 shows that, in terms of bias, there is “no free lunch” for VAR estimation: the worst-case VAR bias is small precisely when the VAR estimator has nearly the same variance as LP. While the worst-case bias can be reduced by increasing the VAR estimation lag length $p$, the proposition shows that this can only happen at the expense of increasing the variance. If we include so many lags that the worst-case bias is zero (cf. Corollary 3.2), then the VAR estimator must necessarily be asymptotically equivalent with LP.
@@
For future reference we briefly discuss how the worst-case mean squared error (MSE) of the VAR estimator depends on the imposed bound on misspecification. Based on Propositions 3.1 and 3.2 as well as Corollary A.2, we define the asymptotic MSE of the VAR and LP estimators as follows:
@@
Impose Assumption 2.1 and $\zeta=1/2$. Then
@@
Impose Assumption 2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\beta}_{h})>\operatorname{aVar}(\hat{\delta}_{h})$. Consider the model-averaging estimator $\hat{\theta}_{h}(\omega)\equiv\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$, and denote its asymptotic MSE by $\operatorname{aMSE}(\hat{\theta}_{h}(\omega))$. Then
@@ 4.2 Worst-case coverage
Impose Assumption 2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\delta}_{h})>0$. Then
@@
This is an immediate consequence of Corollaries 3.1 and 4.1. ∎
Based on this corollary, Figure 4.1 provides a complete characterization of the robustness-efficiency trade-off for VAR confidence intervals. It plots the worst-case coverage probability as a function of the ratio of standard errors for VAR and LP, given significance level $a=10\%$ and different values of $M$. The shaded area depicts an empirically relevant range of standard error ratios obtained in four empirical applications from Ramey (2016). [^6] We replicate Ramey’s identification schemes for monetary, tax, government spending, and technology shocks. The shaded area shows the 10th to 90th percentiles of standard error ratios at horizons exceeding 1 year. See LABEL:app:emp_se for details. We see that, even for $M=1$ (corresponding to a noise-to-signal ratio of $1/T$), the worst-case coverage probability is below 48% whenever the asymptotic standard deviation of the VAR estimator is less than half that of LP—a value that is typical in applied work. Further, at the bottom end of the empirically relevant range, the worst-case coverage probability is essentially zero as soon as $M\geq 1$. It is only at the very right side of the figure—when the VAR includes enough lags to remove nearly all bias, thus increasing the standard error almost to that of LP—that the VAR confidence interval has coverage close to the nominal level.
Figure 4.1: Worst-case asymptotic coverage probability of the conventional 90% VAR confidence interval. Horizontal axis: relative asymptotic standard deviation of LP vs. VAR. Different lines: different bounds $M$ on $\|\alpha(L)\|$. Shaded area: empirical 10th–90th percentile range of relative standard errors based on Ramey (2016), see LABEL:app:emp_se. The solid horizontal line marks the nominal coverage probability $1-a=90\%$.
@@
The shape and magnitude of the least favorable misspecification is difficult to rule out generally based on economic theory. The least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha_{\ell,h,M}^{\dagger}L^{\ell}$ for VAR coverage is the same as the least favorable one for bias (i.e., the $\alpha(L)$ that achieves the maximum in Proposition 4.1). Since $\operatorname{aBias}(\hat{\delta}_{h})$ is linear in $\alpha(L)$, the least favorable choice given the constraint $\|\alpha(L)\|\leq M$ follows easily from the Cauchy-Schwarz inequality (see the proof of Proposition 4.2 below):
@@
We note two main properties of the least favorable misspecification. First, the magnitude of the MA coefficients $\alpha_{\ell,h,M}^{\dagger}$ decays exponentially as $\ell\to\infty$. In other words, not only is the overall magnitude of the least favorable model misspecification small (as imposed in the noise-to-signal bound), the MA coefficients at long lags are in fact particularly small. Second, numerical examples shown in Section A.1 suggest that the MA coefficients tend to be largest in magnitude at horizon $h$, displaying either a hump-shaped pattern as a function of $\ell$—consistent with economic theories of adjustment costs or learning—or a single zig-zag pattern—consistent with theories of overshooting or lumpy adjustment. We thus view MA dynamics of the worst-case form as empirically and theoretically relevant types of misspecification. [^7] However, the least favorable MA polynomial derived above need not be of interest to researchers who trust that some equations in their SVAR specification are exactly correctly specified, as this imposes the additional restrictions that some linear combinations of the rows of the MA polynomial $\alpha(L)$ equal zero.
@@
The least favorable misspecification is also difficult to detect statistically. Propositions 3.3 and 4.1 imply that, for $\alpha(L)=\alpha^{\dagger}(L;h,M)$, the asymptotic rejection probability of the Hausman test of correct VAR specification equals $r(M;z_{1-a/2})$. When $M=1$ (corresponding to a noise-to-signal ratio of $1/T$), the odds of the Hausman test failing to reject the misspecification are nearly 3-to-1 at significance level $a=10\%$, since $r(1;z_{0.95})=26\%$. At significance level $a=5\%$, the odds are nearly 5-to-1, since $r(1;z_{0.975})=17\%$. Standard ex post model misspecification tests are thus unlikely to indicate a problem even if the potential for undercoverage is severe.
@@
Impose Assumption 2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\beta}_{h})>\operatorname{aVar}(\hat{\delta}_{h})>0$. Consider the joint event $\mathcal{A}_{T}$ that $\theta_{h,T}\notin\operatorname{CI}(\hat{\delta}_{h})$ and the Hausman test in Proposition 3.3 fails to reject misspecification. Then
@@
See Section B.7. ∎
Figure 4.2: Worst-case asymptotic probability of the joint event that the conventional VAR confidence interval fails to cover the true impulse response and yet the Hausman test fails to reject misspecification. Horizontal axis: relative asymptotic standard deviation of LP vs. VAR. The dotted horizontal line marks the nominal significance level $a=10\%$.
Figure 4.2 plots this worst-case probability for a significance level of $a=10\%$, which by Corollary 4.4 depends only on the ratio $\operatorname{aVar}(\hat{\delta}_{h})/\operatorname{aVar}(\hat{\beta}_{h})$. Under correct specification, the probability of the joint event is equal to $a(1-a)$ ($=9\%$ when $a=10\%$). With misspecification, the joint probability instead exceeds 46% when the asymptotic standard deviation of the VAR estimator is less than half that of the LP estimator. As $\operatorname{aVar}(\hat{\delta}_{h})/\operatorname{aVar}(\hat{\beta}_{h})\to 0$, the worst-case joint probability approaches $1-a$. We thus again see that statistical tests may fail to warn against the potential for severe VAR coverage distortions.

### v3

@@ 3.1 Large-sample distributions and confidence interval coverage
See Section˜B.2. ∎
The convergence rate $T^{-\min\{1/2,\zeta\}}$ of the VAR estimator is weakly slower than the $T^{-\min\{1/2,2\zeta\}}$ rate achieved by LP. This is because the VAR estimator suffers from bias of order $T^{-\zeta}$, while the stochastic terms of order $T^{-1/2}$ are the same as they would be in a correctly specified SVAR($p$) model. [^4] The first stochastic term captures sampling uncertainty in the reduced-form impulse responses $\hat{A}^{h}$, while the second term captures uncertainty in the structural impact response vector $\hat{\nu}$. The VAR bias is only asymptotically negligible if $\zeta>1/2$, a much smaller degree of robustness than shown above for LP. The case $\zeta=1/2$ is of particular interest, as then the bias and standard deviation are of the same asymptotic order (see also Schorfheide, 2005). MA terms of order $T^{-1/2}$ can be detected with asymptotic probability strictly between 0 and 1 by specification tests, as will be shown in Section˜3.2.
The asymptotic bias is due to two forces: first, the coefficient matrix $\hat{A}$ is biased due to the endogeneity caused by the MA terms, and second, the VAR estimator extrapolates the horizon-$h$ impulse response based on a parametric formula $\hat{A}^{h}$ that does not hold exactly in the true VARMA model (2.1). This is more easily seen in the special case of a univariate model $y_{t}=\rho y_{t-1}+[1+T^{-\zeta}\alpha(L)]\varepsilon_{t}$ with $n=m=1$, in which case
@@
where $\hat{\rho}=\hat{A}$ is the AR(1) coefficient from an OLS regression of $y_{t}$ on $y_{t-1}$. [^5] Lag augmentation of the VAR impulse response estimator as in Inoue and Kilian (2020) may reduce the first term in the bias formula, but it does not affect the second term. While the dynamic responses estimated by the VAR are prone to bias, the shock identification per se is not, in the sense that the VAR’s estimated impact (horizon-0) response is identical to the doubly robust LP estimate (see Section˜2.2).
@@
The preceding results imply that the conventional LP confidence interval is robust to misspecification while the conventional VAR interval is not. We define the level-(1a1-a) LP and VAR confidence intervals using the standard formulae:
@@
Here $z_{1-a/2}$ is the 1a/21-a/2 quantile of the standard normal distribution, and $\operatorname{aVar}(\hat{\beta}_{h})$ and $\operatorname{aVar}(\hat{\delta}_{h})$ are the asymptotic variances of the leading (order-$T^{-1/2}$) stochastic terms in the representations of the LP and VAR estimators in Propositions˜3.1 and 3.2; explicit formulae for the asymptotic variances are given in Corollary˜A.2 in Section˜A.3, which also implies that $\operatorname{aVar}(\hat{\beta}_{h})\geq\operatorname{aVar}(\hat{\delta}_{h})$. None of the results below would change if we replaced the asymptotic variances with the conventional consistent estimates of these (that assume correct specification, as implemented in standard econometric software packages). [^6] Under Assumption 2.1, homoskedastic standard errors suffice. For LP, Heteroskedasticity and Autocorrelation Robust inference would generally be required under the weaker assumptions in LABEL:app:hetero, though simple heteroskedasticity-robust standard errors suffice under the assumptions discussed by Montiel Olea and Plagborg-Møller (2021) and Xu (2023).
@@
Under Assumption˜2.1 and $\zeta>1/4$, $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=1-a$. If moreover $\operatorname{aVar}(\hat{\delta}_{h})>0$ and $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$, then $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\delta}_{h}))=\lim_{T\to\infty}\{1-r\left(T^{1/2-\zeta}b_{h};z_{1-a/2}\right)\}$, where $b_{h}\equiv\operatorname{aBias}(\hat{\delta}_{h})/\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}$, $r(b;c)\equiv P_{Z\sim N(0,1)}(|Z+b|>c)=\Phi(-c-b)+\Phi(-c+b)$, and $\Phi(\cdot)$ is the standard normal distribution function.
@@
Considering separately the three cases $\zeta\in(1/4,1/2)$, $\zeta=1/2$, and $\zeta>1/2$, the result is an immediate consequence of Propositions˜3.1 and 3.2. ∎
LP robustly controls coverage when $\zeta>1/4$, while the VAR confidence interval generically has coverage converging to zero for $\zeta\in(1/4,1/2)$, and strictly below the nominal level 1a1-a for $\zeta=1/2$. Intuitively, the VAR confidence interval has the right width (the same as in the correctly specified case) but the wrong location due to the bias.
@@ 3.2 Hausman misspecification test
Impose Assumption˜2.1, $\zeta>1/4$, and $\operatorname{aVar}(\hat{\beta}_{h})>\operatorname{aVar}(\hat{\delta}_{h})>0$. Then the asymptotic rejection probability of the Hausman test equals
@@
where $b_{h}$ and $r(\cdot,\cdot)$ were defined in Corollary˜3.1.
@@
Considering separately the three cases $\zeta\in(1/4,1/2)$, $\zeta=1/2$, and $\zeta>1/2$, the result follows from Propositions˜3.1 and 3.2 as well as Corollary˜A.2 in Section˜A.3. ∎
As claimed previously, the Hausman test is consistent against MA misspecification of order $T^{-\zeta}$ with $\zeta\in(1/4,1/2)$, except in the knife-edge case where $\operatorname{aBias}(\hat{\delta}_{h})=0$. When $\zeta=1/2$ and $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$, the asymptotic rejection probability is strictly between the significance level $a$ and 1. In Section˜4 we will use the Hausman test to quantify the difficulty of detecting especially pernicious types of model misspecification.
3.3 Lag length selection
We now elaborate further on the role of the estimation lag length. We first discuss the properties of LP when the lag length is selected using standard information criteria. We then show that increasing the lag length robustifies VAR inference by rendering it equivalent with LP inference.
Lag length selection for LP.
LABEL:app:info_crit shows that the LP confidence interval maintains correct asymptotic coverage when the lag length $p$ is selected via the Bayesian Information Criterion (BIC), provided that $\zeta\geq 1/2$. Specifically, the BIC should be applied to an auxiliary VAR in the observed data series $\check{y}_{t}$; the selected lag length then determines the number $p$ of lags to control for in subsequent LP inference (which otherwise discards the auxiliary VAR). The resulting LP inference is robust to model selection errors that are known to cause difficulties for VAR inference (Leeb and Pötscher, 2005; Kilian and Lütkepohl, 2017, chapter 2.6.5). At a high level, the greater reliability of LP inference with data-dependent lag length is a consequence of the double robustness discussed earlier, see Belloni and Chernozhukov (2013) and Chernozhukov et al. (2018).
While the BIC suffices in theory for valid local projection inference, we follow Kilian and Lütkepohl (2017) and recommend that researchers employ the more conservative Akaike Information Criterion (AIC) in practice. The reason is that, in finite samples, the downsides of under-specifying the lag length outweigh the slight inefficiency associated with over-selecting the lag length. Section˜5 demonstrates that this lag length selection procedure delivers LP confidence intervals with accurate coverage in realistic DGPs.
Long-lag VARs.
One simple way to remove the asymptotic bias of the VAR estimator is to control for sufficiently many lags—typically many more lags than indicated by conventional information criteria. This is because in this case the estimator is asymptotically equivalent with the LP estimator. See Plagborg-Møller and Wolf (2021) and Xu (2023) for related results in models without explicit MA misspecification.
@@ 3.3 Lag length selection
Suppose the model (2.2) written in companion form (2.1) satisfies Assumption˜2.1 and $\zeta>1/4$. Let $\tilde{\check{y}}_{t}$ denote the stationary solution to equation (2.2) when $\alpha(L)=0$. If $\varepsilon_{j^{*},t-\ell}\in\operatorname{span}(\tilde{\check{y}}_{t-1},\dots,\tilde{\check{y}}_{t-p})$ for all $\ell=1,\dots,h$, then $\operatorname{aBias}(\hat{\delta}_{h})=0$ and $\operatorname{aVar}(\hat{\delta}_{h})=\operatorname{aVar}(\hat{\beta}_{h})$. In particular, these results obtain if either of the following two sufficient conditions hold:
@@
See Section˜B.3. ∎
We see that, the larger the impulse horizon $h$ of interest, the larger is the estimation lag length $p$ required for bias reduction. In fact, Section˜4 shows that the only way to guarantee that the asymptotic bias of the VAR estimator is zero is to control for so many lags that LP and VAR are asymptotically equivalent.
4 VAR inference under bounded misspecification
To show that the fragility of VARs is likely to matter in practice, we now investigate the worst-case properties of VAR procedures under a tight constraint on the amount of misspecification. We prove that the conventional VAR confidence interval is robust if, and only if, LP and VAR intervals coincide asymptotically. VARs with short-to-moderate lag lengths suffer from severe coverage distortions even for small amounts of misspecification that are hard to rule out either economically or statistically. Beyond increasing the lag length, an alternative strategy to fix VAR undercoverage is to use a larger bias-aware critical value; however, we show that the resulting confidence interval is usually wider than the LP interval. Finally, we show that all conclusions extend to the case of joint inference on multiple impulse responses.
@@ 4.1 Worst-case bias and mean-squared error
where we define the norm $\|\alpha(L)\|\equiv\sqrt{\sum_{\ell=1}^{\infty}\operatorname*{trace}\{D\alpha_{\ell}^{\prime}D^{-1}\alpha_{\ell}\}}$. Suppose we are willing to impose a priori that the noise-to-signal ratio is at most $M^{2}/T$ for some constant $M\in(0,\infty)$. For small $M^{2}/T$, this roughly means that a fraction $M^{2}/T$ of the variance of the model’s error term is due to the misspecification. This corresponds to restricting the parameter space for $\alpha(L)$ to all absolutely summable lag polynomials that satisfy $\|\alpha(L)\|\leq M$. In the following we will consider the worst-case properties of the VAR estimator over this parameter space, treating the other (consistently estimable) parameters $(A,H,D)$ as fixed.
@@
Impose Assumption˜2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\delta}_{h})>0$. Then
@@
where we recall the definition $b_{h}=\operatorname{aBias}(\hat{\delta}_{h})/\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}$. Recall also that $\operatorname{aVar}(\hat{\beta}_{h})$ and $\operatorname{aVar}(\hat{\delta}_{h})$ do not depend on $\alpha(L)$.
@@
The claim is a special case of LABEL:thm:bias_general in LABEL:app:joint. ∎
@@
Proposition˜4.1 shows that VAR estimators must trade off efficiency and robustness: the worst-case VAR bias is small precisely when the VAR estimator has nearly the same variance as LP. While the worst-case bias can be reduced by increasing the VAR estimation lag length $p$, the proposition shows that this can only happen at the expense of increasing the variance. If we include so many lags that the worst-case bias is zero (cf. Corollary˜3.2), then the VAR estimator must necessarily be asymptotically equivalent with LP.
@@
For future reference we briefly discuss how the worst-case mean squared error (MSE) of the VAR estimator depends on the imposed bound on misspecification. Based on Propositions˜3.1 and 3.2 as well as Corollary˜A.2, we define the asymptotic MSE of the VAR and LP estimators as follows:
@@
Impose Assumption˜2.1 and $\zeta=1/2$. Then
@@
Impose Assumption˜2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\beta}_{h})>\operatorname{aVar}(\hat{\delta}_{h})$. Consider the model-averaging estimator $\hat{\theta}_{h}(\omega)\equiv\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$, and denote its asymptotic MSE by $\operatorname{aMSE}(\hat{\theta}_{h}(\omega))$. Then
@@ 4.2 Worst-case coverage
Impose Assumption˜2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\delta}_{h})>0$. Then
@@
This is an immediate consequence of Corollaries˜3.1 and 4.1. ∎
Based on this corollary, Figure˜4.1 provides a complete characterization of the robustness-efficiency trade-off for VAR confidence intervals. It plots the worst-case coverage probability as a function of the ratio of standard errors for VAR and LP, given significance level $a=10\%$ and different values of $M$. The shaded area depicts an empirically relevant range of standard error ratios obtained in four empirical applications from Ramey (2016). [^7] We replicate Ramey’s identification schemes for monetary, tax, government spending, and technology shocks. The shaded area shows the 10th to 90th percentiles of standard error ratios at horizons exceeding 1 year. See the online replication materials for details. We see that, even for $M=1$ (corresponding to a noise-to-signal ratio of $1/T$), the worst-case coverage probability is below 48% whenever the asymptotic standard deviation of the VAR estimator is less than half that of LP—a value that is typical in applied work. Further, at the bottom end of the empirically relevant range, the worst-case coverage probability is essentially zero as soon as $M\geq 1$. It is only at the very right side of the figure—when the VAR includes enough lags to remove nearly all bias, thus increasing the standard error almost to that of LP—that the VAR confidence interval has coverage close to the nominal level.
Figure 4.1: Worst-case asymptotic coverage probability of the conventional 90% VAR confidence interval. Horizontal axis: relative asymptotic standard deviation of VAR vs. LP. Different lines: different bounds $M$ on $\|\alpha(L)\|$. Shaded area: empirical 10th–90th percentile range of relative standard errors based on Ramey (2016), see the online replication materials for details. The solid horizontal line marks the nominal coverage probability $1-a=90\%$.
@@
The shape and magnitude of the least favorable misspecification is difficult to rule out generally based on economic theory. The least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha_{\ell,h,M}^{\dagger}L^{\ell}$ for VAR coverage is the same as the least favorable one for bias (i.e., the $\alpha(L)$ that achieves the maximum in Proposition˜4.1). Since $\operatorname{aBias}(\hat{\delta}_{h})$ is linear in $\alpha(L)$, the least favorable choice given the constraint $\|\alpha(L)\|\leq M$ follows from the Cauchy-Schwarz inequality (see the proof of LABEL:thm:bias_general in LABEL:app:joint):
@@
We note two main properties of the least favorable misspecification. First, the magnitude of the MA coefficients $\alpha_{\ell,h,M}^{\dagger}$ decays exponentially as $\ell\to\infty$. In other words, not only is the overall magnitude of the least favorable model misspecification small (as imposed in the noise-to-signal bound), the MA coefficients at long lags are in fact particularly small. Second, numerical examples shown in Section˜A.1 suggest that the MA coefficients tend to be largest in magnitude at horizon $h$, displaying either a hump-shaped pattern as a function of $\ell$—consistent with economic theories of adjustment costs or learning—or a single zig-zag pattern—consistent with theories of overshooting or lumpy adjustment. We thus view MA dynamics of the worst-case form as empirically and theoretically relevant. [^8] However, the least favorable MA polynomial derived above need not be of interest to researchers who trust that some equations in their SVAR specification are exactly correctly specified, as this imposes the additional restrictions that some linear combinations of the rows of the MA polynomial $\alpha(L)$ equal zero.
@@
The least favorable misspecification is also difficult to detect statistically. Propositions˜3.3 and 4.1 imply that, for $\alpha(L)=\alpha^{\dagger}(L;h,M)$, the asymptotic rejection probability of the Hausman test of correct VAR specification equals $r(M;z_{1-a/2})$. When $M=1$ (corresponding to a noise-to-signal ratio of $1/T$), the odds of the Hausman test failing to reject the misspecification are nearly 3-to-1 at significance level $a=10\%$, since $r(1;z_{0.95})=26\%$. At significance level $a=5\%$, the odds are nearly 5-to-1, since $r(1;z_{0.975})=17\%$. Standard ex post model misspecification tests are thus unlikely to indicate a problem even if the potential for undercoverage is severe.
@@
Impose Assumption˜2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\beta}_{h})>\operatorname{aVar}(\hat{\delta}_{h})>0$. Consider the joint event $\mathcal{A}_{T}$ that $\theta_{h,T}\notin\operatorname{CI}(\hat{\delta}_{h})$ and the Hausman test in Proposition˜3.3 fails to reject misspecification. Then
@@
See Section˜B.6. ∎
Figure 4.2: Worst-case asymptotic probability of the joint event that the conventional VAR confidence interval fails to cover the true impulse response and yet the Hausman test fails to reject misspecification. Horizontal axis: relative asymptotic standard deviation of VAR vs. LP. The dotted horizontal line marks the nominal significance level $a=10\%$.
Figure˜4.2 plots this worst-case probability for a significance level of $a=10\%$, which by Corollary˜4.4 depends only on the ratio $\operatorname{aVar}(\hat{\delta}_{h})/\operatorname{aVar}(\hat{\beta}_{h})$. Under correct specification, the probability of the joint event is equal to a(1a)a(1-a) ($9\%$ when $a=10\%$). With misspecification, the joint probability instead exceeds 46% when the asymptotic standard deviation of the VAR estimator is less than half that of the LP estimator. As $\operatorname{aVar}(\hat{\delta}_{h})/\operatorname{aVar}(\hat{\beta}_{h})\to 0$, the worst-case joint probability approaches 1a1-a. We thus again see that statistical tests may fail to warn against the potential for severe VAR coverage distortions.